OpenStudy (anonymous):

One set of values is p = 5, q = 1, and r = 2.Find p when q = 10 and r = 50.@awkwardpanda

OpenStudy (awkwardpanda):

m and n be constants p = mq = 1/nr 5 = m = 1 / 2n m = 5 and n = 1/10 q = 10 p = 5 * 10 = 50 r = 2 p = 1 / 50 * 10 = 1 / 5 . -anyone can correct me if I'm wrong-

OpenStudy (anonymous):

correct